Find th eq. of the locus of the point which moves such that its distance frm x-y+1+0 is twice its distance from x+y+6=0.

Grade:12

1 Answers

Nehal Wani
13 years ago

Let the point be (x,y)  
|x-y+1|/√2=2|x+y+6|/√2 
|x-y+1|=2|x+y+6| 
x-y+1=±2(x+y+6) 
x+3y+11=0 or 3x+y+13=0